# SoK: On Finding Common Ground in Loss Landscapes Using Deep Model Merging Techniques

### v1

@@ V-C Heuristic Weightings
AdaMerging [97, 98] builds upon Task Arithmetic, learning merging coefficients for task vectors adaptively by minimizing prediction entropy over unlabeled test data. The idea of minimizing prediction entropy originates in test-time adaptation schemes [99, 100] and has been related to finding large, flat basins associated with enhanced generalization capacity [67]. AdaMerging demonstrates better multi-task performance than any other merging method in this section, approaching that of an ensemble. However, it still falls short of the accuracy of a monolithic multi-task model.
@@
Other approaches analyze model activations to determine the importance of particular network components to the final merged model. RegMean [21] merges models trained on separate datasets by computing weight matrices that are the least squares fit to their activation patterns. That is, it attempts to find a weight matrix in each layer that, when presented with an input from the dataset $X_{i}$ corresponding to the source model $\theta_{i}$, produces activations with minimal $L^{2}$ distance from those of $\theta_{i}$. Rather than attempting to match activation maps, ZipIt! [101] seeks to merge only redundant model components whose activations are highly correlated and preserves task-specific components. Intuitively, this prevents destructive interference between task-specific parameters. ZipIt! demonstrates strong performance in multi-task settings—even when models are trained independently.
@@
Some applications require the training of neural networks on data held by edge computing devices without moving data to other locations [102, 103, 104, 105]. A common paradigm for such contexts is Federated Learning (FL) where models are trained separately on edge devices and then merged at a central server using FedSGD [27], a simple average over model parameters. Federated Averaging (FedAvg) [27, 106] generalizes this technique to deal with non-IID data distributions, which degrade the efficacy of simple averaging. FedAvg weights the model parameters returned by each device by the number of training data samples used to update the model parameters, see Eq. (missing) 5. Empirically, this heuristic handles non-IID training well [27]. Additionally, under IID distributions, FedAvg is equivalent to FedSGD. FedAvg is defined formally as:
$\theta_{t+1}^{\star}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{t}^{k}$ (5)
@@
V-D Information Theoretic Weightings
@@
Fisher Weighted Averaging [25] uses the Fisher information of model parameters as weights for a weighted average of parameters. Due to the massive size of the Fisher matrix for neural networks, they use a diagonal approximation to make this method tractable. MaTS [26] improves this method by producing a block-diagonal approximation of the Fisher matrix via Kronecker Factorization [107, 108, 109] and solving the resulting linear system of equations for the optimal merged parameter values using the Conjugate Gradient method [110].
V-E Limitations of Weight Aggregation
@@
Weight aggregation methods have similar advantages to ensembles, such as enhanced generalization, but with a significantly reduced cost. However, they are only applicable and well-defined for merging models of the same architecture. This property precludes their use for more general forms of knowledge transfer than those studied in this section. Another drawback is that their performance degrades as models diverge in their training trajectory. The studies surveyed generally attempt to merge fine-tuned models from the same pre-trained checkpoint; attempts to perform weight aggregation naively on independently trained models are usually unsuccessful [22]. Similarly, weight aggregation between models trained on tasks with significantly different data distributions tends to be unsuccessful. An exception is ZipIt! [101], which gracefully handles divergence in training trajectory and merging between models with different layer sizes—this is because it only merges redundant neurons in each layer. ZipIt! also circumvents issues regarding neuron alignment between layers, as they do not naively merge neurons.
@@
The tendency for weight aggregation methods to fail when training conditions differ may be explained by underlying permutation misalignment between neural networks [47, 55] or by significant differences in the underlying learned features [59, 22], which make it infeasible to directly combine corresponding parameters. Still, the constraints necessary for weight aggregation to succeed illustrate interesting properties of model training behavior. The ability to interpolate smoothly between fine-tuned models originating from the same checkpoint without increases in loss suggests that those models lie in the same linearly connected region. The predictable success of interpolating between fine-tuned models aligns with empirical data indicating that at a certain point in a model’s training trajectory, the final linearly connected objective basin to which they will converge is pre-determined—even if the training distribution is changed after this point in the trajectory [47].
@@
A linear path between weights has also been shown to reflect a linear path between learned features in some scenarios [39], particularly under the pre-training and fine-tuning paradigm [111]. This result may explain why some aggregation procedures might succeed without accounting for permutation invariance. Moreover, the success of weight steering methods indicates that—at least within an objective basin—we can accurately assign directions in parameter space to task-specific behavior. This property has also been noted by Gueta et al. [82], who observe a clustering in parameter space for LLMs fine-tuned from the same pre-trained checkpoint on semantically similar tasks.
@@
Figure 4: Mode Connectivity: Many objective basins are equivalent up to a permutation in neurons. Models in distinct objective basins can be “transported” close to one another by applying appropriate permutations to their units.
VI Neuron Alignment
@@
As opposed to naive weight aggregation schemes, neuron alignment techniques reduce potential degradation in the resulting merged model by first aligning the neurons of each source model and then performing weight aggregation. This prevents weight aggregation from destroying important features from source models, which may occur if the source models are permuted relative to one another as described in Section III-B2. Neuron alignment schemes outperform naive weight aggregation when models are trained independently or otherwise lie in different regions of the objective landscape, often revealing linear paths between distinct modes.
VI-A Learning Permutation Matrices
@@
Entezari et al. [16] conjecture that all objective modes are linearly connected if one accounts for permutation symmetries between modes. To this end, various neuron alignment techniques attempt to explicitly determine a permutation matrix that can align the neurons of one model to those of another. Ainsworth et al. [19] present two methods for aligning neurons via permutation matrices: activation matching and weight matching. Activation matching aligns each corresponding layer between two source models by finding a permutation matrix that minimizes the $L^{2}$ distance of their activation patterns. Weight matching finds a permutation matrix that attempts to maximize the Frobenius product between weight matrices. Intuitively, weight matching aligns weights that are similar to one another and produce similar activations as a result. These methods are set up as a sum-of-bilinear-assignments problem and solved using an iterative approximation algorithm. Both methods are competitive with their third proposed method, which is to learn a permutation matrix directly via a straight-through-estimator (STE). This method updates the weights of the second source model, $\theta_{b}$, to directly minimize the loss barrier (Eq. (missing) 1), projecting the result to the nearest realizable permutation of $\theta_{b}$’s parameters. Notably, all three methods observe that their performance degrades as model width decreases and similarly worsens for models in the initial stages of training.
@@
Tatro et al. [51] relate the neuron alignment problem to the assignment problem [112] and optimal transport. They align neurons to maximize the correlation between their pre-activations and solve the resulting alignment problem using the Hungarian Algorithm [113]. They also provide theoretical results indicating that such alignment tightens the upper bound on the loss of interpolated models. Interestingly, they also report better performance with wider model architectures. Probabilistic Federated Neural Matching (PFNM) [114] seeks to align the parameters of edge models to a global model in the federated setting using a Bayesian nonparametric model. Federated Learning with Matched Averaging [58] takes inspiration from PFNM and FedAvg by using the Hungarian Algorithm to align corresponding layers, reducing the communication burden associated with merging in the federated setting.
VI-B Learning Alignment Maps
@@
Some methods are not constrained to use permutation matrices but instead seek to learn unconstrained “soft” maps that enable neuron alignment for merging. Optimal Transport Fusion (OTFusion) [20] poses the model merging problem as an optimal transport problem and aligns neurons based on activation and weight similarities using an exact optimal transport solver, requiring only a single pass over all layers as opposed to other iterative neuron alignment methods (see Section VI-A). The optimal transport framework allows for merging between models with differing layer sizes. While previous methods explicitly constrain optimization to return permutation matrices, OTFusion imposes no such constraint. Instead, it finds a Wasserstein barycenter between the distributions encoded by each source model. This amounts to optimization over the Birkhoff polytope [115] in $n$ dimensions whose vertices are the traditional permutation matrices in $S_{n}$ (if the model layer sizes are equivalent). In this regard, OTFusion is a generalization of prior methods that relate neuron alignment to the discrete assignment problem. Once again, we observe that OTFusion’s performance improves with increased model width.
@@
OTFusion has since been extended to handle transformer architectures [116]—specifically dealing with residual connections, multi-head attention, and layer normalization—as well as to align neurons across layers in heterogenous architectures (as opposed to only within corresponding layers) [117]. Despite their good performance, even aligned neural networks sometimes demonstrate high loss barriers on the path between them. Jordan et al. [118] attribute this problem to a reduction in the variance of hidden layer activations during interpolation and propose an intermediate scaling layer that improves the performance of merged models.
VI-C Limitations of Neuron Alignment
@@
Merging procedures combined with neuron alignment seem to resolve issues raised by prior work associated with merging models from independent training trajectories. Permutation-based methods also empirically reinforce Entezari et al.’s conjecture [16] that many, if not all, objective modes are linearly connected if one accounts for permutation symmetries. One drawback of the current state-of-the-art for neuron alignment is that these methods are often forced to perform greedy layer-wise alignment for tractability when, in principle, permutation symmetries may exist across layers as well. Therefore, there may be a set of permutation invariances that these methods do not account for—as well as other yet undiscovered forms of misalignment between neurons.
@@
Additionally, the issue of merging models trained on different tasks remains. STE [19] and AdaMerge [97] both show promise in this regard but require data samples and backpropagation—purely manipulating weight matrices seems insufficient. Yamada et al. [22] explore this problem empirically demonstrating that merging performance degrades as the training distributions of source models diverge. They conjecture that the growing differences in loss geometry that arise as datasets diverge require optimization over samples from a mixed-task dataset. They use Dataset Condensation [119] to cheaply construct mixed datasets and optimize merged parameters by using the STE method, with promising results (albeit only on MNIST and Fashion-MNIST, both known to be relatively “easy” datasets [47].
@@
Ito et al. [59] investigate the success of neuron alignment methods further by analyzing the net effect of neuron alignment methods on the properties of weight matrices. Proponents of permutation-based methods often paint an intuitive picture of models being transported into the same convex objective basin, evoking an image of reduced $L^{2}$ distance between weights. However, their investigation shows a poor correlation between neuron alignment and $L^{2}$ distance. Rather, the benefit of neuron alignment seems to be predicated on the alignment of dominant singular directions in the weight matrices of each layer. They also show that models independently trained on similar tasks tend to show a similar distribution of singular values. This suggests that difficulties merging models trained independently on the same task may be explained by differences in their singular directions rather than by fundamental differences in learned features. Jin et al. [21] indicate the converse: that models stemming from the same pre-trained model (i.e., that lie in the same objective basin) show no evidence of permutation between weights. This implies that permutation-based procedures may only be necessary when source models are trained independently.
@@
These studies further indicate that differences in learned features between models trained on different tasks may render procedures based purely on weight manipulation ineffective. Merging across tasks may require optimization via backpropagation or avoiding merging task-specific components altogether as in ZipIt! [101].
VII Insights into Training
@@
VII-A Loss Landscape Macrophenomena
@@
Mode Convexity: Objective basins are locally convex, permitting one to travel a short distance and still encounter a meaningful neural network solution with similar behavior. More importantly, models with roughly equivalent or lower loss but meaningfully different prediction behavior exist in proximity to any given solution. Fast Geometric Ensembling (FGE) [50] implicitly takes advantage of this phenomenon to rapidly discover quality ensemble members after training only a single model and Stochastic Weight Averaging (SWA) [64] exploits this phenomenon by averaging rather than ensembling several models from the same training trajectory. These studies suggest that local minima in the objective landscape contain models with sufficiently diverse behavior to gain a performance advantage by ensembling or merging them—despite their proximity in parameter space.
@@
Figure 5: Mode Determinism: As a training trajectory proceeds from the initialization point (pictured as the red circle here) its path may diverge to various minima in the loss landscape due to SGD noise (e.g., batch order) or changes in the underlying data distribution, but after a certain number of steps the trajectory of a model reaches a stable point (pictured as the red diamond). Any further fine-tuning from this point onwards, even on slightly different data distributions, will result in models lying in the same convex objective basin (red stars).

### v2

@@ 4.3. Data-Aware Weightings
AdaMerging (133; 139) builds upon Task Arithmetic, learning merging coefficients for task vectors adaptively by minimizing prediction entropy over unlabeled test data. The idea of minimizing prediction entropy originates in test-time adaptation schemes (119; 82) and has been related to finding large, flat basins associated with enhanced generalization capacity (17). AdaMerging demonstrates better multi-task performance than any other merging method in this section, approaching that of an ensemble. However, it still falls short of the accuracy of a monolithic multi-task model.
@@
Other approaches analyze model activations to determine the importance of particular network components to the final merged model. RegMean (64) merges models trained on separate datasets by computing weight matrices that are the least squares fit to their activation patterns. That is, it attempts to find a weight matrix in each layer that, when presented with an input from the dataset $X_{i}$ corresponding to the source model $\theta_{i}$, produces activations with minimal $L^{2}$ distance from those of $\theta_{i}$. Rather than attempting to match activation maps, ZipIt! (108) seeks to merge only redundant model components whose activations are highly correlated and preserves task-specific components. Intuitively, this prevents destructive interference between task-specific parameters. ZipIt! demonstrates strong performance in multi-task settings—even when models are trained independently. Evolutionary Model Merging is a similar technique that uses an evolutionary algorithm to learn a policy for whether to compose layers from each source model in sequence or to learn merging weights for each parameter in a given layer (4).
@@
Some applications require the training of neural networks on data held by edge computing devices without moving data to other locations (127; 57; 56; 26). A common paradigm for such contexts is Federated Learning (FL) where models are trained separately on edge devices and then merged at a central server using FedSGD (91), a simple average over model parameters. Federated Averaging (FedAvg) (91; 109) generalizes this technique to deal with non-IID data distributions, which degrade the efficacy of simple averaging. FedAvg weights the model parameters returned by each device by the number of training data samples used to update the model parameters, see Eq. (missing) 5. Empirically, this heuristic handles non-IID training well (91). Under IID distributions, FedAvg is equivalent to FedSGD. FedAvg is defined formally as:
(5) $\theta_{t+1}^{\star}=\sum_{k=1}^{K}\frac{n_{k}}{n}\theta_{t}^{k}$
@@
4.4. Information Theoretic Weightings
@@
Fisher Weighted Averaging (90) uses the Fisher information of model parameters as weights for a weighted average of parameters. Due to the massive size of the Fisher matrix for neural networks, they use a diagonal approximation to make this method tractable. MaTS (110) improves this method by producing a block-diagonal approximation of the Fisher matrix via Kronecker Factorization (88; 43; 7) and solving the resulting linear system of equations for the optimal merged parameter values using the Conjugate Gradient method (47).
4.5. Limitations of Weight Aggregation
@@
Weight aggregation methods have similar advantages to ensembles, such as enhanced generalization, but with a significantly reduced cost. However, they are only applicable and well-defined for merging models of the same architecture. This property precludes their use for more general forms of knowledge transfer than those studied in this section. Another drawback is that their performance degrades as models diverge in their training trajectory. The studies surveyed generally attempt to merge fine-tuned models from the same pre-trained checkpoint; attempts to perform weight aggregation naively on independently trained models are usually unsuccessful (131). Similarly, weight aggregation between models trained on tasks with significantly different data distributions tends to be unsuccessful. An exception is ZipIt! (108), which gracefully handles divergence in training trajectory and merging between models with different layer sizes—this is because it only merges redundant neurons in each layer. ZipIt! also circumvents issues regarding neuron alignment between layers, as they do not naively merge neurons.
@@
The tendency for weight aggregation methods to fail when training conditions differ may be explained by underlying permutation misalignment between neural networks (40; 106) or by significant differences in the underlying learned features (61; 131), which make it infeasible to directly combine corresponding parameters. Still, the constraints necessary for weight aggregation to succeed illustrate interesting properties of model training behavior. The ability to interpolate smoothly between fine-tuned models originating from the same checkpoint without significant increases in loss along this path suggests that those models lie in the same linearly connected region, this notion of Mode Convexity is illustrated in Fig. 3. The predictable success of interpolating between fine-tuned models aligns with empirical data indicating that at a certain point in a model’s training trajectory, the final linearly connected objective basin to which they will converge is pre-determined—even if the training distribution is changed after this point in the trajectory (40).
@@
A linear path between weights has also been shown to reflect a linear path between learned features in some scenarios (142), particularly under the pre-training and fine-tuning paradigm (140). This result may explain why some aggregation procedures might succeed without accounting for permutation invariance. Moreover, the success of weight steering methods indicates that—at least within an objective basin—we can accurately assign directions in parameter space to task-specific behavior. This property has also been noted by Gueta et al. (44), who observe a clustering in parameter space for LLMs fine-tuned from the same pre-trained checkpoint on semantically similar tasks.
@@
Figure 5. Mode Connectivity: Many objective basins are equivalent up to a permutation in neurons. Models in distinct objective basins can be “transported” close to one another by applying appropriate permutations to their units.
5. Neuron Alignment
Table 3. Overview of Neuron Alignment techniques
Technique Finds Permutation Matrix Finds “Soft” Mapping Requires Data Requires Labels
Weight Matching (3)
Heuristic Neuron Alignment (115)
Probabilistic Federated Neural Matching (135)
FedMA (120)
OTFusion (107)
Transformer OTFusion (60)
@@ 5. Neuron Alignment
As opposed to naive weight aggregation schemes, neuron alignment techniques reduce potential degradation in the resulting merged model by first aligning the neurons of each source model and then performing weight aggregation. This prevents weight aggregation from destroying important features from source models, which may occur if the source models are permuted relative to one another as described in Section 2.2.2. Neuron alignment schemes outperform naive weight aggregation when models are trained independently or otherwise lie in different regions of the objective landscape, often revealing linear paths between distinct modes. Table 3 lists the works we review in this section along with some of their high level features.
5.1. Learning Permutation Matrices
@@
Entezari et al. (30) conjecture that all objective modes are linearly connected if one accounts for permutation symmetries between modes, this notion of Mode Connectivity is illustrated in Fig. 5. To this end, various neuron alignment techniques attempt to explicitly determine a permutation matrix that can align the neurons of one model to those of another. Ainsworth et al. (3) present two methods for aligning neurons via permutation matrices: activation matching and weight matching. Activation matching aligns each corresponding layer between two source models by finding a permutation matrix that minimizes the $L^{2}$ distance of their activation patterns. Weight matching finds a permutation matrix that attempts to maximize the Frobenius product between weight matrices. Intuitively, weight matching aligns weights that are similar to one another and produce similar activations as a result. These methods are set up as a sum-of-bilinear-assignments problem and solved using an iterative approximation algorithm. Both methods are competitive with their third proposed method, which is to learn a permutation matrix directly via a straight-through-estimator (STE). This method updates the weights of the second source model, $\theta_{b}$, to directly minimize the loss barrier (Eq. (missing) 1), projecting the result to the nearest realizable permutation of $\theta_{b}$’s parameters. Notably, the performance of all three methods degrades as model width decreases and similarly worsens for models in the initial stages of training.
@@
Tatro et al. (115) relate the neuron alignment problem to the assignment problem (14) and optimal transport. They align neurons to maximize the correlation between their pre-activations and solve the resulting alignment problem using the Hungarian Algorithm (71). They also provide theoretical results indicating that such alignment tightens the upper bound on the loss of interpolated models. Interestingly, this method also performs better on wider model architectures. Probabilistic Federated Neural Matching (PFNM) (135) seeks to align the parameters of edge models to a global model in the federated setting using a Bayesian nonparametric model. Federated Learning with Matched Averaging (120) takes inspiration from PFNM and FedAvg by using the Hungarian Algorithm to align corresponding layers, reducing the communication burden associated with merging in the federated setting.
5.2. Learning Alignment Maps
@@
Rather than constraining themselves to learning permutation matrices, some methods instead seek to learn unconstrained “soft” maps that enable neuron alignment for merging. Optimal Transport Fusion (OTFusion) (107) poses the model merging problem as an optimal transport problem and aligns neurons based on activation and weight similarities using an exact optimal transport solver, requiring only a single pass over all layers as opposed to the other iterative neuron alignment methods presented in Section 5.1. The optimal transport framework allows for merging between models with differing layer sizes. While previous methods explicitly constrain optimization to return permutation matrices, OTFusion imposes no such constraint. Instead, it finds a Wasserstein barycenter between the distributions encoded by each source model. This amounts to optimization over the Birkhoff polytope (143) in $n$ dimensions whose vertices are the traditional permutation matrices in $S_{n}$ (if the model layer sizes are equivalent). In this regard, OTFusion is a generalization of prior methods that relate neuron alignment to the discrete assignment problem. Once again, we note that OTFusion’s performance improves with increased model width.
@@
OTFusion has since been extended to handle transformer architectures (60)—specifically dealing with residual connections, multi-head attention, and layer normalization—as well as to align neurons across layers in heterogenous architectures (as opposed to only within corresponding layers) (97). Despite their good performance, even aligned neural networks sometimes demonstrate high loss barriers on the path between them. Jordan et al. (65) attribute this problem to a reduction in the variance of hidden layer activations during interpolation and propose an intermediate scaling layer that improves the performance of models merged using neuron alignment techniques, reducing the loss barrier between models anywhere from 60-100%.
5.3. Limitations of Neuron Alignment
@@
Merging procedures combined with neuron alignment seem to resolve issues raised by prior work associated with merging models from independent training trajectories. Permutation-based methods also empirically reinforce Entezari et al.’s conjecture (30) that many, if not all, objective modes are linearly connected if one accounts for permutation symmetries. One drawback of the current state-of-the-art for neuron alignment is that these methods are often forced to perform greedy layer-wise alignment for tractability when, in principle, permutation symmetries may exist across layers as well. Therefore, there may be a set of permutation invariances that these methods do not account for—as well as other yet undiscovered forms of misalignment between neurons.
@@
Additionally, the issue of merging models trained on different tasks remains. STE (3) and AdaMerge (133) both show promise in this regard but require data samples and backpropagation—purely manipulating weight matrices seems insufficient. Yamada et al. (131) explore this problem, empirically demonstrating that merging performance degrades as the training distributions of source models diverge. They conjecture that the growing differences in loss geometry that arise as datasets diverge require optimization over samples from a mixed-task dataset. They use Dataset Condensation (137) to cheaply construct mixed datasets and optimize merged parameters by using the STE method, with promising results (albeit only on MNIST and Fashion-MNIST, both known to be relatively “easy” datasets (40)).
@@
Ito et al. (61) investigate the success of neuron alignment methods further by analyzing the net effect of neuron alignment methods on the properties of weight matrices. Proponents of permutation-based methods often paint an intuitive picture of models being transported into the same convex objective basin, evoking an image of reduced $L^{2}$ distance between weights. However, their investigation shows a poor correlation between neuron alignment and $L^{2}$ distance. Instead, the benefit of neuron alignment seems to be predicated on the alignment of dominant singular directions in the weight matrices of each layer. They also show that models independently trained on similar tasks tend to show a similar distribution of singular values. This suggests that difficulties in merging models trained independently on the same task may be explained by differences in their singular directions rather than by fundamental differences in learned features. Jin et al. (64) indicate the converse: that models stemming from the same pre-trained model (i.e., that lie in the same objective basin) show no evidence of permutation between weights. This implies that permutation-based procedures are only necessary when source models are trained independently. Ito et al. (61) also explain the correlation between greater merging success and model width by the fact that as model width increases, the distribution of dominant singular values becomes more concentrated. The existence of fewer dominant feature directions to align makes the alignment process more manageable (see Section 6.2.1).
@@
These studies further indicate that differences in learned features between models trained on different tasks may render procedures based purely on weight manipulation ineffective. Merging across tasks may require optimization via backpropagation or avoiding merging task-specific components altogether, as in ZipIt! (108).
6. Insights into Training
@@
6.1. Loss Landscape Macrophenomena
@@ 6.1.1. Mode Convexity:
Objective basins are locally convex, permitting one to travel a short distance and still encounter a meaningful neural network solution with similar behavior. More importantly, models with roughly equivalent or lower loss but meaningfully different prediction behavior exist in proximity to any given solution. Fast Geometric Ensembling (FGE) (41) implicitly takes advantage of this phenomenon to rapidly discover quality ensemble members after training only a single model and Stochastic Weight Averaging (SWA) (62) exploits this phenomenon by averaging rather than ensembling several models from the same training trajectory. These studies suggest that local minima in the objective landscape contain models with sufficiently diverse behavior to gain a performance advantage by ensembling or merging them—despite their proximity in parameter space.
@@
Figure 6. Mode Determinism: As a training trajectory proceeds from the initialization point (pictured as the red circle here) its path may diverge to various minima in the loss landscape due to SGD noise (e.g., batch order) or changes in the underlying data distribution, but after a certain number of steps the trajectory of a model reaches a stable point (pictured as the red diamond). Any further fine-tuning from this point onwards, even on slightly different data distributions, will result in models lying in the same convex objective basin (red stars).
6.1.2. Mode Determinism: